OpenStudy (anonymous):

What is the measure of each base angle of an isosceles triangle if its vertex angle measures 38 degrees and its 2 congruent sides measure 23 units?

OpenStudy (anonymous):

 What is the measure of each base angle of an isosceles triangle if its vertex angle measures 38 degrees and its 2 congruent sides measure 23 units?

OpenStudy (anonymous):

the sum of the angles of atrianlge os 180 degrees use this simple formula to find the measure of the base angle. . base angle= (180-38)/2 =71 degrees.

OpenStudy (anonymous):

why i divide it by 2?

OpenStudy (anonymous):

i got 71 degrees. but how can each angle is to 71 degrees if trainlges equal to 180?

OpenStudy (anonymous):

because, the base angle of an isosceles triangle are equal.. let's check A+B+C=180 71+71+38= 180 180 =180

OpenStudy (anonymous):

is it clear to you?

OpenStudy (anonymous):

yes thank you.